Frances is 2 years older than Anna. The sum of their ages is twenty-six. How old is Frances?
Answer by JulietG(1812)

F + A = 26
F = A + 2
Substitute the value of F from the second equation into the first.
(A+2) + A = 26
2A + 2 = 26
Subtract 2 from each side.
2A = 24
Divide each side by 2
A = 12
If Anna is 12, then Frances is 14.